# Quantum-Assisted Vehicle Routing: Realizing QAOA-based Approach on Gate-Based Quantum Computer

### v2

@@ Abstract.
The Vehicle Routing Problem (VRP) is a fundamental combinatorial optimization challenge with broad applications in logistics and transportation. In this work, we present a quantum-assisted framework that integrates the Quantum Approximate Optimization Algorithm (QAOA) with a link-based formulation of VRP. Our approach encodes flow conservation and subtour elimination directly into the cost Hamiltonian, preserving graph structure while avoiding ancillary qubits. We design and implement the full pipeline on a gate-based quantum computer, including problem encoding, circuit synthesis, and execution on IBM Quantum System One. Experimental results on small VRP instances highlight the effects of penalty scaling, coefficient normalization, and circuit depth on solution feasibility under hardware noise. While scalability remains constrained by circuit complexity and decoherence, the study demonstrates a practical pathway for implementing VRP on quantum hardware and identifies methodological directions for advancing near-term quantum optimization.
@@
journal: JATSjournalvolume: 37journalnumber: 4article: 111publicationmonth: 11ccs: Computer systems organization Quantum computingccs: General and reference Experimentationccs: Mathematics of computing Combinatorial optimizationcopyright: acmlicensedjournal: TQCjournalyear: 2025journalvolume: 1journalnumber: 1article: 1publicationmonth: 1doi: 10.1145/0000003
@@ 1. Introduction
Figure 1. Solving the VRP using QAOA involves a formulation of the link based constrained minimization form to the unconstrained quadratic binary representation, which is then cast as an Ising Hamiltonian required for QAOA. The high level circuit is transpiled into one and two qubit basic gates for implementation on the IBM-Rensselaer quantum computer. The circuit is run repeatedly to optimize the variational parameters, and a final sampling provides the solution. The problem graph shown here is for illustrative purposes only. Section IV describes the specific problem studied.
@@
In this work, we go beyond a direct application of QAOA to VRP by co-designing quantum circuits with the problem’s graph structure. We develop both the algorithmic formulation and its implementation framework on gate-based quantum hardware, i.e., IBM Quantum System One. While the problem is NP-hard, the inherent graph structure of VRP makes it a compelling candidate for quantum approaches. Unlike approaches that precalculate the clusters on a classical computer (Feld et al., 2019), or use time expanded versions of the VRP to trade off qubits for constraint simplification (Palackal et al., 2023), we design the quantum solution based on a link based optimization formulation. Using directional links as decision variables we naturally capture the graph structure of VRP and this allows us to incorporate subtour elimination constraints and flow conservation directly into the cost Hamiltonian while avoiding ancillary qubits.
To the best of our knowledge, while there have been studies on using quantum simulation for VRP, this paper is among the first efforts to design a comprehensive pipeline from modeling to implementing VRP on a gate-based quantum computer. Moreover, the study examines key methodological aspects, including problem encoding, circuit depth, parameter optimization, and hardware related limitations such as noise and decoherence. We found that the choice of penalty scaling and coefficient normalization significantly affects solution feasibility in QAOA. In particular, normalization of Hamiltonian terms improved performance on larger VRP instances, while penalty values around twice the sum of edge weights provided consistently better feasibility. Together, these analyses shed light on the present challenges and point to promising directions for advancing quantum optimization techniques in real-world routing applications.
@@ 2. Literature Review
Special-purpose quantum computers such as quantum annealers are built to exploit quantum mechanics for specific problem domains. Similar to thermal simulated annealing (Kirkpatrick et al., 1983), quantum annealing is a process that finds the lowest-energy configuration of a system by gradually transforming from a simple quantum state to one representing the optimal solution (Finnila et al., 1994). It was shown to outperform classical methods for the traveling salesperson problem (Martoňák et al., 2004), and has since found commercial use through companies such as D-Wave that are now offering devices with over 5000 qubits which see use in logistics, finance, and machine learning (Boothby et al., 2020). However, special-purpose computers exhibit a speedup that is problem-dependent, and can only perform a narrow set of tasks and lack the flexibility to run any general quantum algorithm (Yarkoni et al., 2022). The capacitated vehicle routing problem has been approached by a hybrid algorithm utilizing quantum annealing, with the results showing no clear benefit over classical solvers (Feld et al., 2019).
@@
While interest in quantum computing for intelligent transportation systems has been expressed only recently (Burkacky et al., 2020)(Zhuang et al., 2024), the combinatorial optimization problems of TSP and VRP, have received significant attention from a quantum computing perspective (Martoňák et al., 2004)(Leonidas et al., 2023). Classical methods exist to solve the problem exactly, for example branch and bound (Theurich et al., 2021), which scale exponentially with problem size. On the other hand metaheuristic algorithms can provide approximate solutions, like Tabu search (Sicilia et al., 2016) and genetic algorithms (Baker and Ayechew, 2003), but they are not guaranteed to find the exact optimal solution in large problems. QAOA was applied to small scale TSP instances, giving high quality solutions at limited depth with a mixer Hamiltonian corresponding to the problem constraints (Ruan et al., 2020). The performance of QAOA simulation on a VRP formulation without subtour elimination constraints was found to depend strongly on the classical optimization method and the depth of the ansatz since training time can increase exponentially, but this may be prevented by a good initialization (Azad et al., 2022). Another study discovered that QAOA biases solutions toward lower energies but fails to distinguish between feasible solutions of different costs, limiting its effectiveness for VRP (Fitzek et al., 2024). The authors conjecture that this may be due to a small energy gap in the Hamiltonian representation, which could be mitigated by reformulating the problem to increase energy separation. A detailed comparison of different VRP formulations and quantum algorithms demonstrated that QAOA with constrained optimization by linear approximation (COBYLA) delivered the highest probability of sampling the optimal configuration given enough post optimization sampling shots (Harwood et al., 2021).
@@ 3.1. Quantum Computing
The state of multiple qubits lies in an exponential state space. If there are $n$ qubits, each of the $2^{n}$ basis states has a complex coefficient, $\alpha_{i}$, that dictates the probability, $|\alpha_{i}|^{2}$, of measuring that basis state. It is for this reason that quantum computing cannot be efficiently simulated on classical computers (Bremner et al., 2011). For example, 50 qubits will require $2^{50}$ complex numbers to describe the state, and any circuit operations will require multiplication with a matrix of size $2^{50}\times 2^{50}$. If each coefficient was represented with just one byte, the state vector would take over 1000 Terabytes to store in a classical computer.
@@ 3.3. Vehicle Routing Problem
where $S$ is any subset of the node set $\mathcal{N}$, consisting of $n$ total nodes. Meaning that any selected subset of the nodes must have at least one active edge to some node in the complement set. Without these explicit constraints, subtours will appear in the solution space which seemingly satisfy all the other constraints and may have lower cost than the real solution. For example, in an 8-node, 2-vehicle problem there could be the set of routes $0\rightarrow 1\rightarrow 2\rightarrow 0$, $0\rightarrow 3\rightarrow 4\rightarrow 0$, and a subtour $5\rightarrow 6\rightarrow 7\rightarrow 5$. This set satisfies all constraints except subtour elimination.
@@
where $c_{r}$ is the cost of a route, $\delta_{i,r}$ is 1 if route $r$ visits node $i$, and $\mathcal{R}$ is the set of routes. While this formulation has fewer constraints, the number of routes in a fully connected network is $|\mathcal{N}|!$, which have to be generated explicitly in a preprocessing step before any optimization algorithm is run.
Another encoding method uses the time step expansion by including time steps into the decision variable $x^{k}_{v.t}$, which is 1 if the $k$th vehicle visits the node $v$ at time step $t$. The optimization problem is then
@@ 4.1. QAOA Formulation
A randomly instantiated VRP with 3 nodes and 2 vehicles was formulated using the link based formulation in Equation 9 as a combinatorial optimization problem with a linear cost function and linear constraints. Qiskit optimization libraries were used to create the instance of the problem and convert to QUBO and the Ising Hamiltonian. While this package is no longer officially supported by IBM, the functionality may be adapted to work with the latest version of Qiskit.
@@
The formulation was converted to a quadratic unconstrained binary optimization (QUBO) problem by creating an augmented objective function similar to the Lagrange multiplier approach (Glover et al., 2018). This method adds to the original cost function quadratic penalties corresponding to each constraint, resulting in an unconstrained problem which has the same global optimum as the original problem. For example, $C_{0}$ can be converted to a penalty term $P(1-x_{1,0}-x_{1,2}+2x_{1,0}x_{1,2})$, where $P$ is a sufficiently large positive scalar, and $C_{6}$ may be added as $Px_{1,2}x_{2,1}$. Note that these transformations are unique to binary variables.
@@ 4.3. Results and Interpretation
The results obtained from running the hybrid QAOA algorithm for the VRP provide insights into the effectiveness and scalability of quantum computing for combinatorial optimization tasks. The final sampling using the trained parameters resulted in a distribution of results, shown in Figure 6. It can be seen that the most likely bit-string is 111010, which corresponds to $x_{0,1},x_{0,2},x_{1,0},x_{2,0}$ having the value 1, and the remaining variables are 0.
@@
Our findings also highlight the rapid growth in quantum gate counts as the problem scales. Larger instances that require more depth quickly become too large to run practically. For example a 4 node problem with depth of 2 generates a 14 qubit circuit consisting of 3207 $R_{Z}$, 1642 $R_{X}$, 928 echoed cross-resonance, and 214 Pauli-X gates. This problem took over 2 hours to run on the quantum computer and generated an infeasible solution. In fact, the top 10 most sampled states were all infeasible. Similarly, a 5 node problem with depth 2 transpiled into a 30 qubit circuit of about 30,000 basic gates. The number of circuit gates increases linearly with $p$ as plotted in Figure 8(a), while the total gates increases exponentially against the number of nodes in the problem. However, a more representative measure of circuit depth from a noise accumulation perspective is the maximum number of two qubit gates in a path from the start of the circuit to the end. This metric exhibits an exponential increase as shown in Figure 8(b).
@@ 4.4. Circuit complexity
The main complexity in the case of the VRP arises from the subtour elimination constraints, which increase exponentially with the number of nodes, since there is one constraint for every subset of nodes and $O(2^{n})$ subsets. Each constraint with more than 3 variables can be broken into multiple smaller constraints by introducing new binary variables, which is the approach used for generating the QUBO. As a result, there ends up being an exponentially increasing number of terms in the QUBO and, therefore, the Ising Hamiltonian. To illustrate, the VRP was instantiated for a number of nodes, and the number of variables in the QUBO were counted to plot the graph in Figure 9.
@@ 4.5. Penalty Parameter Tuning and Normalization Effects
With normalization applied, QAOA for the 4-node problem was able to recover the correct solution among the most frequently sampled bitstrings. Without normalization, however, the algorithm failed to consistently identify the optimal route in the high probability region of the output distribution.
@@ 4.6. Limitations and Future Considerations
While the QAOA has demonstrated potential for solving combinatorial optimization problems, as the problem size increases, it encounters several challenges. The observed exponential growth in circuit complexity caused by the problem mapping leads to significant decoherence and noise accumulation, reducing solution accuracy. Without constraints, the number of qubits would increase with the square of the number of nodes in the VRP because the problem requires representation of pairwise interactions between nodes. However, the inclusion of subtour elimination constraints adds an exponentially increasing number of slack variables and Pauli terms to the problem Hamiltonian. While methods such as unbalanced penalization (Montañez-Barrera et al., 2024) can reduce the number of variables, they introduce additional interaction terms into the Hamiltonian, which may not be a favorable tradeoff given the overhead of swap and other two qubit gates required to implement those interactions on physical quantum hardware. Moreover, the penalty coefficients have to be optimized classically for every problem beforehand.

### v3

@@ Abstract.
The Vehicle Routing Problem (VRP) is a fundamental combinatorial optimization challenge with broad applications in logistics and transportation. In this work, we present a quantum-assisted framework that integrates the Quantum Approximate Optimization Algorithm (QAOA) with a link-based formulation of VRP. Our approach encodes flow conservation and subtour elimination directly into the cost Hamiltonian, preserving graph structure while minimizing resource requirements for practical hardware implementation. We design and implement the full pipeline on a gate-based quantum computer, including problem formulation, encoding, circuit synthesis, and execution on IBM Quantum System One. Experimental results on small VRP instances highlight the effects of penalty scaling, coefficient normalization, and circuit depth on solution feasibility under hardware noise. While scalability remains constrained by circuit complexity and decoherence, the study demonstrates a practical pathway for implementing VRP on quantum hardware and identifies methodological directions for advancing near-term quantum optimization.
@@
copyright: noneccs: Computer systems organization Quantum computingccs: General and reference Experimentationccs: Mathematics of computing Combinatorial optimization
@@ 1. Introduction
Figure 1. Solving the VRP using QAOA involves a formulation of the link based constrained minimization form to the unconstrained quadratic binary representation, which is then cast as an Ising Hamiltonian required for QAOA. The high level circuit is transpiled into one and two qubit basic gates for implementation on the IBM-Rensselaer quantum computer. The circuit is run repeatedly to optimize the variational parameters, and a final sampling provides the solution. The problem graph shown here is for illustrative purposes only. Section 4 describes a specific example problem.
@@
This work presents a hardware grounded study of QAOA for the Vehicle Routing Problem (VRP), with novelty arising from the co-design of problem formulation, circuit construction, and experimental evaluation on real quantum hardware rather than simulation-only analysis. The main contributions are:
  • We develop a directional edge-based formulation that preserves the graph structure of VRP and embeds flow conservation and subtour elimination directly into the cost Hamiltonian. Unlike time-expanded or route-order encodings, this approach avoids explicit route sequencing variables and reduces qubit and gate overhead, making it more suitable for near-term devices.
  • We implement the full workflow, from VRP encoding and QUBO construction to Ising mapping, circuit synthesis, transpilation, and execution on IBM Quantum System One. This provides empirical insight into QAOA behavior under realistic noise, connectivity, and depth constraints, which are not captured by idealized simulations.
  • We quantitatively evaluate how VRP formulations impact qubit count and two-qubit gate depth after hardware-aware transpilation, showing that circuit depth is a primary bottleneck for constrained routing problems on current superconducting architectures.
  • We demonstrate that both penalty parameter selection and coefficient normalization critically influence solution feasibility. In particular, penalties on the order of twice the sum of edge weights and normalization of Hamiltonian coefficients significantly improve feasible-solution sampling on larger instances.
Overall, the novelty of this work lies in its practical, hardware executed evaluation of QAOA for VRP, offering concrete design insights that clarify current limitations and inform future quantum optimization studies in the NISQ regime.
@@ 2. Literature Review
Special-purpose quantum computers such as quantum annealers are built to exploit quantum mechanics for specific problem domains. Similar to thermal simulated annealing (Kirkpatrick et al., 1983), quantum annealing is a process that finds the lowest-energy configuration of a system by gradually transforming from a simple quantum state to one representing the optimal solution (Finnila et al., 1994). It was shown to outperform classical methods for the traveling salesperson problem (Martoňák et al., 2004), and has since found commercial use through companies such as D-Wave that are now offering devices with over 5000 qubits which see use in logistics, finance, and machine learning (Boothby et al., 2020). However, special-purpose computers exhibit a speedup that is problem-dependent, and can only perform a narrow set of tasks and lack the flexibility to run any general quantum algorithm (Yarkoni et al., 2022). The capacitated vehicle routing problem has been approached by a hybrid algorithm utilizing quantum annealing, with the results showing no clear benefit over classical solvers (Feld et al., 2019). This method used a cluster-then-route approach which loses global connections, and subsequent routing cannot be guaranteed to give the true optimum.
@@
While interest in quantum computing for intelligent transportation systems has been expressed only recently (Burkacky et al., 2020)(Zhuang et al., 2024), the combinatorial optimization problems of TSP and VRP, have received significant attention from a quantum computing perspective (Martoňák et al., 2004)(Leonidas et al., 2023). Classical methods exist to solve the problem exactly, for example branch and bound (Theurich et al., 2021), which scale exponentially with problem size. On the other hand metaheuristic algorithms can provide approximate solutions, like Tabu search (Sicilia et al., 2016) and genetic algorithms (Baker and Ayechew, 2003), but they are not guaranteed to find the exact optimal solution in large problems. QAOA was applied to small scale TSP instances, giving high quality solutions at limited depth with a mixer Hamiltonian corresponding to the problem constraints (Ruan et al., 2020). The performance of QAOA simulation on a VRP formulation without subtour elimination constraints was found to depend strongly on the classical optimization method and the depth of the ansatz since training time can increase exponentially, but this may be prevented by a good initialization (Azad et al., 2022). However, the authors acknowledge the appearance of subtours in the QAOA results and recommend explicit penalization. Another study discovered that QAOA biases solutions toward lower energies but fails to distinguish between feasible solutions of different costs, limiting its effectiveness for VRP (Fitzek et al., 2024). The authors conjecture that this may be due to a small energy gap in the Hamiltonian representation, which could be mitigated by reformulating the problem to increase energy separation. A detailed comparison of different VRP formulations and quantum algorithms demonstrated that QAOA with constrained optimization by linear approximation (COBYLA) delivered the highest probability of sampling the optimal configuration given enough post optimization sampling shots (Harwood et al., 2021).
@@ 3.1. Quantum Computing
The state of multiple qubits lies in an exponential state space. If there are $n$ qubits, each of the $2^{n}$ basis states has a complex coefficient, $\alpha_{i}$, that dictates its probability, $|\alpha_{i}|^{2}$, of measuring that basis state. It is for this reason that quantum computing cannot be efficiently simulated on classical computers (Bremner et al., 2011). For example, 50 qubits will require $2^{50}$ complex numbers to describe the state, and any circuit operations will require multiplication with a matrix of size $2^{50}\times 2^{50}$. If each coefficient was represented with just one byte, the state vector would take over 1000 Terabytes to store in a classical computer.
@@ 3.3. Vehicle Routing Problem
where $S$ is any subset of the node set $\mathcal{N}$, consisting of $n$ total nodes. Meaning that any selected subset of the nodes must have at least one active edge to some node in the complement set. Without these explicit constraints, subtours will appear in the solution space which seemingly satisfy all the other constraints and may have lower cost than the real solution. For example, in an 8-node, 2-vehicle problem there could be the set of routes $0\rightarrow 1\rightarrow 2\rightarrow 0$, $0\rightarrow 3\rightarrow 4\rightarrow 0$, and a subtour $5\rightarrow 6\rightarrow 7\rightarrow 5$. This set satisfies all constraints except subtour elimination. This issue has been noted in previous works such as (Azad et al., 2022), which saw infeasible solutions appear in simulated QAOA output as their formulation did not include penalties for subtours.
@@
where $c_{r}$ is the cost of a route, $\delta_{i,r}$ is 1 if route $r$ visits node $i$, and $\mathcal{R}$ is the set of routes. While this formulation has fewer constraints, the number of routes in a fully connected network is $|\mathcal{N}|!$, which have to be generated explicitly in a preprocessing step before any optimization algorithm is run. Since each route corresponds to a qubit in this scheme, it would require a factorial number of qubits to implement making it infeasible except for very small problems.
Another encoding method uses the time step expansion by including time steps into the decision variable $x^{k}_{v.t}$, which is $1$ if the $k$th vehicle visits the node $v$ at time step $t$. This is also known as a sequence-based formulation (Harwood et al., 2021). The optimization problem is then
@@
A similar formulation is adopted in (Fitzek et al., 2024) by introducing a routing variable that tracks the order in which each node is visited, eliminating the need to consider subtour constraints. In the Results section we compare the resource requirements for these formulations. For now, we continue with the link (edge)-based formulation discussed first.
@@ 4.1. QAOA Formulation
A randomly instantiated VRP with 3 nodes and 2 vehicles was formulated using the link based formulation in Equation 9 as a combinatorial optimization problem with a linear cost function and linear constraints. Qiskit optimization libraries were used to create the instance of the problem and convert to QUBO and the Ising Hamiltonian. While this package is no longer officially supported by IBM, the functionality may be adapted to work with the latest version of Qiskit. The coefficients of the following instance are the squared euclidean distance between the nodes indicated by the subscripts.
@@
The formulation was converted to a quadratic unconstrained binary optimization (QUBO) problem by creating an augmented objective function similar to the Lagrange multiplier approach (Glover et al., 2018). This method adds to the original cost function quadratic penalties corresponding to each constraint, resulting in an unconstrained problem which has the same global optimum as the original problem. For example, $C_{0}$ can be converted to a penalty term $P(1-x_{1,0}-x_{1,2}+2x_{1,0}x_{1,2})$, where $P$ is a sufficiently large positive scalar, and $C_{6}$ may be added as $Px_{1,2}x_{2,1}$. Note that these transformations are unique to binary variables. Inequality constraints involving more than two binary variables typically require the introduction of auxiliary slack variables to be represented exactly as quadratic penalty terms.
@@ 4.3. Results and Interpretation
The results obtained from running the hybrid QAOA algorithm for the VRP provide insights into the effectiveness and scalability of quantum computing for combinatorial optimization tasks. The final sampling using the trained parameters resulted in a distribution of results, shown in Figure 6. It can be seen that the most likely bit-string is 111010, which corresponds to $x_{0,1},x_{0,2},x_{1,0},x_{2,0}$ having the value 1, and the remaining variables are 0. It has been verified with a classical solver that this is the optimum solution.
@@
Our findings also highlight the rapid growth in quantum gate counts as the problem scales. Larger instances that require more depth quickly become too large to run practically. For example a 4 node problem with depth of 2 generates a 14 qubit circuit consisting of 3207 $R_{Z}$, 1642 $R_{X}$, 928 echoed cross-resonance, and 214 Pauli-X gates. This problem took over 2 hours to run on the quantum computer and generated an infeasible solution. In fact, the top 10 most sampled states were all infeasible. Similarly, a 5 node problem with depth 2 transpiled into a 30 qubit circuit of about 30,000 basic gates. The number of circuit gates increases linearly with $p$ as plotted in Figure 8(a), while the total gates increases exponentially against the number of nodes in the problem. However, a more representative measure of circuit depth from a noise accumulation perspective is the maximum number of two qubit gates in a path from the start of the circuit to the end, shown in Figure 8(b). Using leave-one-out cross-validation on four problem sizes, a quadratic fit exhibits lower prediction error than an exponential one, indicating that the observed growth is consistent with quadratic scaling over the tested range.
@@ 4.4. Circuit complexity
Using the edge-based formulation, the main complexity in the case of the VRP arises from the subtour elimination constraints, which increase exponentially with the number of nodes, since there is one constraint for every subset of nodes and $O(2^{n})$ subsets. Each constraint with more than 3 variables can be broken into multiple smaller constraints by introducing new binary variables, which is the approach used for generating the QUBO. As a result, there ends up being an exponentially increasing number of terms in the QUBO and, therefore, the Ising Hamiltonian. To illustrate, the VRP was instantiated for a number of nodes, and the number of variables in the QUBO were counted to plot the graph in Figure 9.
@@
4.4.1. Choice of Routing Formulation
Among the three routing formulations examined earlier, we selected the edge-based formulation with subtour elimination for quantum implementation. Although time-expanded and order-based encodings remove subtours structurally, their substantially larger QUBO sizes and resulting two-qubit gate depths render them impractical on current quantum hardware.
The first issue is that of the number of qubits. The IBM-Rensselaer quantum computer has 127 qubits, which puts a limit on the size of any problem that can be executed on it. Figure 10(a) shows the number of qubits needed to run the formulations discussed in Section 3.3, highlighting the multiplicative effect of introducing routing variables for each vehicle. The time-expanded and route-order formulations require the same number of binary variables, therefore their graphs overlap. Note that we do not consider vehicle-flow formulations as they are more relevant when demands and capacities are considered (Irnich et al., 2014). For example, the two-index flow formulation from (Munari et al., 2016) produces a QUBO requiring more variables than any of the other methods.
Figure 10(b) shows the two-qubit gate depth for the same formulations after transpilation to our target hardware. For four vehicles there are two isolated points corresponding to the time-expanded and route-ordered formulations in the 5-node case, since there cannot be four vehicles in a smaller problem, and for larger problems the number of qubits required is greater than the quantum computer has available, so transpilation fails. The edge-based formulation requires a lower two-qubit gate depth across all cases.
Consequently, by both practical measures the edge-based formulation provides the most feasible way to execute QAOA for problems in this regime on available hardware.
(a) Number of qubits vs. problem size for different number of vehicles
(b) Two-qubit gate depth vs. problem size for different number of vehicles
Figure 10. Comparison of qubits and two-qubit gate depth for edge-based vs. time-expanded (Palackal et al., 2023) and route-order (Fitzek et al., 2024) formulations. $p=2$ for all cases.
@@ 4.5. Penalty Parameter Tuning and Normalization Effects
With normalization applied, QAOA for the 4-node problem was able to recover the correct solution among the most frequently sampled bitstrings. Without normalization, however, the algorithm failed to consistently identify the optimal route in the high probability region of the output distribution. This observation was consistent across multiple randomly generated 4 and 5 node instances.
@@
Beyond demonstrating a single successful instance, the results provide actionable guidance on penalty selection, Hamiltonian normalization, and formulation choice that directly inform how constrained optimization problems should be encoded for near-term quantum processors.
@@ 4.6. Limitations and Future Considerations
While the QAOA has demonstrated potential for solving combinatorial optimization problems, as the problem size increases, it encounters several challenges. The observed exponential growth in circuit complexity caused by the problem mapping leads to significant decoherence and noise accumulation, reducing solution accuracy. Without constraints, the number of qubits would increase with the square of the number of nodes in the VRP because the problem requires representation of pairwise interactions between nodes. However, the inclusion of subtour elimination constraints adds an exponentially increasing number of slack variables and Pauli terms to the problem Hamiltonian. While methods such as unbalanced penalization (Montañez-Barrera et al., 2024) can reduce the number of variables, they introduce additional interaction terms into the Hamiltonian, which may not be a favorable tradeoff given the overhead of swap and other two-qubit gates required to implement those interactions on physical quantum hardware. Moreover, the penalty coefficients would have to be optimized classically for every problem beforehand. Future quantum computers with more qubits and lower two-qubit error rates could take advantage of route-order formulations as they provide sub-exponential scaling, but this becomes favorable only for larger problems.